# Geometry of quantum states and chaos-integrability transition

### v1

@@ III.2 Solutions of the geodesic equations
Here $\lambda$ is a parameter which parametrises the geodesic path. For the metric components given in (24), these equations are given by
@@
where for convenience, we have denoted $A=(N-1)/2$. Since $g_{ab}\dot{x}^{a}\dot{x}^{b}=K^{2}$ is also constant along the geodesic, expanding this constraint and subsequently eliminating $\dot{\phi}(\lambda)$ using eq. (28), we get the following first-order equation for $r(\lambda)$,
@@
where (ϕ0)±subscriptsubscriptitalic-$(\phi_{0})_{\pm}$ denotes the initial values of the $\phi$ coordinates for two solutions. We can fix this constant in such a way that $\phi_{\pm}(\lambda=0)=0$. We also assume that the constant $L\geq 0$, so that the relation $\dot{\phi}(\lambda)=\frac{L(r^{2}+1)}{Ar^{2}}$ implies that both the solutions grows with $\lambda$.
@@ IV Correlation function of GUE matrices and the fidelity susceptibility
As we have discussed in section II.1, the components of the ensemble-averaged QMT are closely related to the two-point connected correlation functions of the relevant operators. To understand this connection more closely in the context of random matrix Hamiltonians, in this section, we compute the correlation function of the form $\braket{\mathcal{H}(t)\mathcal{H}}$ that appears in (14), with the time evolution generated by a Hamiltonian of the type in (1). Then we shall show how one can recover the expressions of the FS from this correlation function, thereby illuminating the important role played by the correlation between the energy levels in determining the behaviour of the geometry of the eigenstates.
In this section, we first evaluate the correlation function by scaling the total Hamiltonian [^10] The scaling is performed to make the connection with an alternative method of evaluating the correlation function presented in Appendix D manifest. Towards the end of this section we provide the expression for the correlation function appearing in (14) for the non-scaled Hamiltonian. as $\tilde{H}=H/\alpha$, where $\alpha=\sqrt{(r^{2}+1)}\sigma$, where $\sigma^{2}$ is the variance of the matrices $H_{0}$ and $\mathcal{H}$, both drawn from independent GUE. Hence, the Hamiltonian we consider is, [^11] In terms of the notation used in section III, =~$\mathcal{H}=\tilde{\mathcal{H}}_{1}$. To keep the discussion simple, here we assume that the total Hamiltonian depends on only one parameter $r$. Therefore, the FS one would get here is the same as the $rr$ component of the QMT in (24).
@@
Thus, the random variables $X$ and $\tilde{H}$ are independent of each other. It can be straightforwardly checked that, similar to $\tilde{H}$, the average variance of the elements $X$ is unity, i.e, $\mathbb{E}\big{(}|X_{ij}|^{2}\big{)}=1$. th Next, we perform these variable changes in the ensemble averaging, namely, $\{H_{0},\mathcal{H}\}\rightarrow\{\tilde{H},X\}$. Hence, the correlation function under consideration becomes,
@@
where $C_{1}=\sigma/\sqrt{r^{2}+1}$ and $C_{2}=(r\sigma)/\sqrt{r^{2}+1}$. Note that the constant factors coming from the Jacobians of the transformations cancel from the corresponding factors in the denominator. Using the fact that the mean and variance of both the variables ($X$ and $\tilde{H}$) are 00 and $1$, respectively, it is easy to check that $\braket{X\tilde{H}}=0$ and $\braket{\tilde{H}^{2}}=N$. Finally, the first term in the second line of the expression can be checked to be related to the spectral form factor by the relation, $\braket{X(t)X}=\frac{\tilde{\mathcal{R}}(t)}{N}$, where we have defined the SFF as
~$\tilde{\mathcal{R}}(t)=Z^{-1}\int\big{|}\text{Tr}\big{(}e^{-i\tilde{H}t}\big{)}\big{|}^{2}~{}P(\tilde{H})~{}\text{d}\tilde{H}~{}.$ (41)
@@
Figure 1: Plot of the correlation function $\tilde{G}(t)=\braket{\mathcal{H}(t)\mathcal{H}}$ (in the Log-Log scale), where the time evolution is generated by the rescaled Hamiltonian (35). Here $N=1000$, $r=0.43$, and the plot shows the result with an average over 100 independent realisations of the Hamiltonian $\tilde{H}$. The red curve denotes the numerical result, while $\tilde{G}(t)$ computed from the formulas in (42) and (97), are respectively shown by the grey, black dashed curves. The plot in the inset shows, more closely, the behaviour of the correlation function at late times (with the same colour coding) and the difference between these two analytical expressions. This difference would be negligible in the limit $N\rightarrow\infty$.
@@ IV.1 Fidelity susceptibility from correlation function
where $\mathcal{R}(t)$ now denotes the SFF associated with $H(r)$, i.e.,
@@
To obtain the expression for the FS, we first note that, from the definition of SFF, we have,
@@
here we have used the fact that $N^{-1}\mathbb{E}\sum_{n}|\braket{n}{\mathcal{H}}{n}|^{2}=(1+Nr^{2})/(N(r^{2}+1))$ in (12). A straightforward computation shows that we get back the expression for the FS,
@@ V Geometric Properties of Quantum metric tensor for integrability breaking Hamiltonian
As we have discussed in the Introduction, one of the main focuses of this paper is to study the geometry of quantum energy eigenstates of systems that show chaos-to-integrability transitions. For this purpose, in this section, we begin by considering the eigenstate space geometry of a Hamiltonian which can be written in the following way,
which is of the same form as the one considered in the previous section (eq.(21)), however, here the matrix $H_{0}$ is taken to be a diagonal matrix whose elements are drawn independently from Gaussian distributions with mean zero and unit variance. This Hamiltonian can be thought of as a generalisation of the standard RP model [33].
@@ V.1 QMT of a $2\times 2$ Integrability breaking Hamiltonian
In this spirit, a two-parameter family of $2\times 2$ random matrices, where two GUE matrices are added to a random diagonal integrable matrix, was used in [42] to obtain an exact analytical expression for the components of the ensemble-averaged QMT. Specifically, the Hamiltonian is of a similar structure as that of the one in (21), where $H_{0}$ is a diagonal matrix whose elements are independent and identically distributed Gaussian random variables. Below, we use this QMT to study its various geometric properties, such as the scalar curvature invariants and the structure of the geodesics in this geometry.
The line element corresponding to the ensemble-averaged QMT obtained from a $2\times 2$ matrix representation of the integrability-breaking Hamiltonian in (54) is given in the $\{r,\phi\}$ coordinates by the following expression [^12] We use a tilde to distinguish the new quantities for the metric in (55) from the analogous quantities for the random matrix case discussed in the previous section. However, to avoid cluttering the notation, we have removed the bar indicating an ensemble average. [42]
@@
Therefore, as one moves from integrability by increasing the value of $r$, the value of the Ricci scalar increases and subsequently attains a constant positive value of $4$, consistent with the discussion of the previous section. Also note that as one decreases $r$, the Ricci scalar monotonically decreases, and in the limit $r\rightarrow 0$, it attains a constant positive value $8/\pi^{2}$. The Ricci scalar is regular for all admissible values of the parameter $r$, indicating that there is no curvature singularity in the geometry. We have shown a plot of the Ricci scalar in Fig. 2.
@@ V.2 Solutions of the geodesic equations
We now proceed to obtain the geodesic curves for the metric in (55). We first notice that, for the metric in (55) as well, $\phi$ is a cyclic coordinate. Hence, we can obtain the conserved quantity associated with the Killing vector $K^{a}=(0,1)$ to be  [^13] To distinguish from the conserved quantity $L$ of the previous section, here we have denoted the conserved quantity associated with the cyclic coordinate $\phi$ to $\tilde{L}$.
@@
We first provide numerical solutions to the above equation for two cases, when the $\dot{\theta}(\lambda)$ is positive and when it’s negative. Then we will analyse an approximate analytical solution of this equation, which, as we shall show, provides a good approximation of the numerical results.
@@
In Fig. 5 we have shown a comparison of the approximate expression (64) with the numerically obtained solution. As can be seen, the solution in eq. (64) actually provides quite a good approximation to the numerical solution for the entire range of $\theta$. For comparison, we have also plotted a simple linear expression $\theta(\lambda)\approx A_{1}(c-\lambda)$, with $c$ being a constant, which shows that close to $\theta=0$ (far away from the integrable phase), the geodesic solution $\theta(\lambda)$ linearly goes to zero.
@@
As can be seen by expanding this expression, for small values of the parameter $\lambda$, $r(\lambda)$ grows linearly, after which quadratic growth takes over. This expression, along with $r(\lambda)$ obtained using the numerical solution for $\theta(\lambda)$ is plotted in Fig. 6 for two different initial values of $r_{0}$. The approximate expression in (65) is in good agreement with the numerical solution for both the initial conditions. Furthermore, from this solution for $r(\lambda)$ we once again see that one can reach any arbitrary value of $r$, for a finite value of the parameter $\lambda$ following this geodesic. Therefore, any point arbitrarily far away from the integrable phase can be reached with a finite distance, i.e., a finite value of the ’complexity’.
@@
Before concluding this section we note that here we have performed an analysis of different geometric quantities for the generalised version of the conventional RP model with $N=2$. While it would be very much desirable to perform a similar analysis for a generic $N$, analytic expressions for the QMT components for an arbitrary value of $N$ are not known. In [42], these components were obtained numerically, and it was shown that the metric components follow different scaling properties in different phases of this model. E.g., when $r\ll r^{*}=1/\sqrt{N}$ (corresponding to the localised phase of the model), the metric components follow a $1/r$ scaling, while in the ergodic phase, having $r\gg 1$, the metric components approach those of the GUE random matrix expressions in (24). Finally, in the non-ergodic extended phase ($r^{*}\ll r\ll 1$), the components $\tilde{g}_{rr}$ and $\tilde{g}_{\phi\phi}/r^{2}$ goes to constants scaling as $N$.
@@ VI Geometry of quantum states for Gaussian $\beta$-ensembles
In this section, and the next, we consider a generalization of the integrability-breaking Hamiltonian considered in the previous section, such that the Dyson index ($\beta$) can take a range of real positive values, not just those of classical Gaussian ensembles. Specifically, in this section, we consider the tridiagonal matrix representation of the so-called Gaussian $\beta$-ensembles, found in [40], a Hamiltonian belonging to the ensemble takes a real symmetric tridiagonal $N\times N$ matrix whose diagonal elements (denoted as $a_{n}$, $n=0,\cdots N-1$) are are independent Gaussian random variables with mean 00, and variance as that of the ensemble, while the sub-diagonal elements (which we call $b_{n}$, with $n=1,\cdots N-1$) are independently distributed random variables with the following distribution
@@
A natural property of the ensemble of Hamiltonians with general $\beta$ is that, by controlling the value of the Dyson index, one can control the strength of the repulsion between the energy levels. In the limit $\beta\rightarrow 0$, the energy levels can be arbitrarily close to each other without any level repulsion, and therefore the level spacing statistics is Poissonian, associated with integrable systems. Increasing the value of $\beta$ from this value, one gradually gets ensembles with different degrees of level repulsion, including those of the classical Gaussian ensembles (with $\beta=1,2,4$), as well as ensembles having intermediate values of $\beta$. Therefore, the Gaussian $\beta$-ensembles can be thought to be showing a chaos-to-integrability transition as $\beta$ is decreased below the GUE value ($\beta=1$) towards zero. Similar to the RP model, apart from the chaos to integrability transition, the $\beta$-ensembles also show Anderson localisation transition for a specific value of the Dyson index. Rewriting $\beta=N^{\gamma}$ for some real parameter $\gamma$, it has been shown recently that $\beta$-ensembles have three different phases, respectively, an ergodic phase ($\gamma\leq 0$), a nonergodic extended phase (for $10<\gamma<$ 1) and a localised phase when $\gamma\geq 1$ [69].
@@
For these ensemble with generic Dyson index, one can follow two different procedures to move to the integrable phase: 1. first, as was the case in section III, we can add a term proportional to random matrix drawn from an ensemble with a fixed general value of the Dyson index $\beta>1$ to a diagonal matrix having independent random variables as the entry. As one takes the proportionality constant (denoted as $r$) to zero the total Hamiltonian makes a transition from chaotic to integrable phase as $r\rightarrow 0$. 2. Secondly, we can consider a Hamiltonian which is a sum of two terms each of which taken from these generalised $\beta$-ensembles and have, in general, different Dyson indices ($\beta_{1}$ and $\beta_{2}$). Then, as one takes either or both of these $\beta_{i}$ close zero and take $r\rightarrow 0$, the total Hamiltonian again makes a transition from chaotic to integrable phase.
[…]

### v2

@@ III.2 Solutions of the geodesic equations
Here $\lambda$ is a parameter which parametrises the geodesic path, and $\Gamma^{a}_{bc}$ denotes the Christoffel symbols which can be computed from the metric components using the standard formulas [61]. For the metric components given in (27), these two equations reduce to
@@
where for convenience, we have denoted A=(N1)/2A=(N-1)/2. Since $g_{ab}\dot{x}^{a}\dot{x}^{b}=K^{2}$ is also constant along the geodesic, expanding this constraint and subsequently eliminating $\dot{\phi}(\lambda)$ using eq. (31), we get the following first-order equation for $r(\lambda)$,
@@
where $(\phi_{0})_{\pm}$ denotes the initial values of the $\phi$ coordinates for two solutions. We can fix this constant in such a way that $\phi_{\pm}(\lambda=0)=0$. We also assume that the constant $L\geq 0$, so that the relation $\dot{\phi}(\lambda)=\frac{L(r^{2}+1)}{Ar^{2}}$ implies that both the solutions grows with $\lambda$.
@@ IV Correlation function of GUE matrices and the fidelity susceptibility
As we have discussed in section II.1, the FS is related to the two-point connected correlation functions of the operator $\mathcal{H}$. To understand this connection more closely in the context of random matrix Hamiltonians, in this section, we compute the correlation function of the form $\braket{\mathcal{H}(t)\mathcal{H}}$ that appears in (16), with the time evolution generated by a Hamiltonian of the type in (1). Then we shall show how one can recover the expressions of the FS from this correlation function, thereby illuminating the important role played by the correlation between the energy levels in determining the behaviour of the FS, and hence, the geometry of the eigenstates.
To begin with, we first evaluate the required correlation function by scaling the total Hamiltonian [^12] The scaling is performed to make the connection with an alternative method of evaluating the correlation function presented in Appendix D manifest. Towards the end of this section we provide the expression for the correlation function appearing in (16) for the non-scaled Hamiltonian. as $\tilde{H}=H/\alpha$, where $\alpha=\sqrt{(r^{2}+1)}\sigma$, where $\sigma^{2}$ is the variance of the matrices $H_{0}$ and $\mathcal{H}$, both drawn from independent GUE. Hence, the Hamiltonian we consider is, [^13] In terms of the notation used in section III, $\mathcal{H}=\tilde{\mathcal{H}}_{1}$. To keep the discussion simple, here we assume that the total Hamiltonian depends on only one parameter $r$. Therefore, the FS one would get here is the same as the $rr$ component of the QMT in (27).
@@
Thus, the random variables $X$ and $\tilde{H}$ are independent of each other. It can be straightforwardly checked that, similar to $\tilde{H}$, the average variance of the elements $X$ is unity, i.e, $\mathbb{E}\big{(}|X_{ij}|^{2}\big{)}=1$. Next, we perform these variable changes in the ensemble averaging, namely, use the transformation, $\{H_{0},\mathcal{H}\}\rightarrow\{\tilde{H},X\}$. Hence, the correlation function under consideration becomes,
@@
where $C_{1}=\sigma/\sqrt{r^{2}+1}$ and $C_{2}=(r\sigma)/\sqrt{r^{2}+1}$. Note that the constant factors coming from the Jacobians of the transformations cancel from the corresponding factors in the denominator $Z$. Using the fact that the mean and variance of both the variables ($X$ and $\tilde{H}$) are 0 and $1$, respectively, it is easy to check that $\braket{X\tilde{H}}=0$ and $\braket{\tilde{H}^{2}}=N$. Finally, the first term in the second line of the expression can be checked to be related to the spectral form factor by the relation, $\braket{X(t)X}=\frac{\tilde{\mathcal{R}}(t)}{N}$, where we have defined the SFF as
$\tilde{\mathcal{R}}(t)=Z^{-1}\int\big{|}\text{Tr}\big{(}e^{-i\tilde{H}t}\big{)}\big{|}^{2}~P(\tilde{H})~\text{d}\tilde{H}~.$ (44)
@@
Figure 1: Plot of the correlation function $\tilde{G}(t)=\braket{\mathcal{H}(t)\mathcal{H}}$ (in the Log-Log scale), where the time evolution is generated by the rescaled Hamiltonian (38). Here $N=1000$, $r=0.43$, and the plot shows the result with an average over 100 independent realisations of the Hamiltonian $\tilde{H}$. The red curve denotes the numerical result, while $\tilde{G}(t)$ computed from the formulas in (45) and (110), are respectively shown by the grey, black dashed curves. The plot in the inset shows, more closely, the behaviour of the correlation function at late times (with the same colour coding) and the difference between these two analytical expressions. This difference would vanish in the limit $N\rightarrow\infty$.
@@ IV.1 Fidelity susceptibility from correlation function
where $\mathcal{R}(t)$ now denotes the SFF associated with a scaled Hamiltonian $H^{\prime}(r)$, i.e.,
@@
To obtain the expression for the FS, we first notice that, from the definition of SFF, we have,
@@
here we have used the fact that $N^{-1}\mathbb{E}\sum_{n}|\braket{n|\mathcal{H}|n}|^{2}=(1+Nr^{2})/(N(r^{2}+1))$ in (14). A straightforward computation shows that we get back the expression for the FS,
@@ V Geometric Properties of Quantum metric tensor for integrability breaking Hamiltonian
As we have discussed in the Introduction, one of the main focuses of this paper is to study the geometry of energy eigenstates of systems that show chaos-to-integrability transitions. For this purpose, in this section, we begin by considering the eigenstate space geometry of a Hamiltonian which can be written in the following way,
$H(r,\phi)=H_{0}+r\cos\phi~\mathcal{H}_{1}+r\sin\phi~\mathcal{H}_{2}~.$ (57)
This Hamiltonian is of the same form as the one considered in the previous section (eq.(24)), however, here the matrix $H_{0}$ is taken to be a diagonal matrix whose elements are drawn independently from Gaussian distributions with mean zero and unit variance. The level spacing statistics of this matrix is therefore given by the Poisson statistics, and hence it is an integrable Hamiltonian [72, 51, 52]. This two-parameter family of Hamiltonians can be thought of as a generalisation of the standard version of the RP model [35].
@@ V.1 QMT of a $2\times 2$ Integrability breaking Hamiltonian
In this spirit, a two-parameter family of $2\times 2$ random matrices, where two GUE matrices are added to a random diagonal integrable matrix, was used in [44] to obtain an exact analytical expression for the components of the ensemble-averaged QMT. Specifically, the Hamiltonian is of a similar structure to that of the one in (24), where $H_{0}$ is a diagonal matrix whose elements are independent and identically distributed Gaussian random variables. Below, we use this QMT to study its various geometric properties, such as the scalar curvature invariants and the structure of the geodesics in this geometry.
The line element corresponding to the ensemble-averaged QMT (see eq. (23)) obtained from a $2\times 2$ matrix representation of the integrability-breaking Hamiltonian in (57) is given in the $\{r,\phi\}$ coordinates by the following expression [^14] We use a tilde to distinguish the new quantities for the metric in (58) from the analogous quantities for the random matrix case discussed in the previous section. However, to avoid cluttering the notation, we have removed the bar indicating an ensemble average. [44]
@@
Therefore, as one moves from integrability by increasing the value of $r$, the value of the Ricci scalar increases and subsequently attains a constant positive value of $4$, consistent with the discussion of the previous section. Also note that as one decreases $r$, the Ricci scalar monotonically decreases, and in the limit $r\rightarrow 0$, it attains a constant positive value $8/\pi^{2}$. The Ricci scalar is regular for all admissible values of the parameter $r$, indicating that there is no curvature singularity in the geometry. We have shown a plot of the Ricci scalar in Fig. 2 to illustrate the features described above.
@@ V.2 Solutions of the geodesic equations
We now proceed to obtain the geodesic curves for the metric in (58). The strategy we follow is very similar to the one outlined in III. Since the parameter space is two-dimensional here as well, the geodesic equations are given by (28), with the Christoffel symbols calculated from the metric components in (58). Once again, our goal is to reduce the second-order equations to first-order ones, which can be integrated subsequently to obtain the geodesic curves. To this end, we first notice that, for the metric in (58) as well, $\phi$ is a cyclic coordinate. Hence, we can obtain the conserved quantity associated with the Killing vector $K^{a}=(0,1)$ to be  [^15] To distinguish from the conserved quantity $L$ of the previous section, here we have denoted the conserved quantity associated with the cyclic coordinate $\phi$ to $\tilde{L}$.
@@
We first provide numerical solutions to the above equation for two cases, when the $\dot{\theta}(\lambda)$ is positive and when it’s negative. Then we will present an approximate analytical solution of this equation, which, as we shall show, provides a good approximation of the numerical results.
@@
In Fig. 5, we have shown a comparison of the approximate expression (67) with the numerically obtained solution. As can be seen, the solution in eq. (67) actually provides quite a good approximation to the numerical solution for the entire range of $\theta$. For comparison, we have also plotted a simple linear expression $\theta(\lambda)\approx A_{1}(c-\lambda)$, with $c$ being a constant, which shows that close to $\theta=0$ (far away from the integrable phase), the geodesic solution $\theta(\lambda)$ linearly goes to zero.
@@
As can be seen by expanding this expression, for small values of the parameter $\lambda$, $r(\lambda)$ grows linearly, after which quadratic growth takes over. This expression, along with $r(\lambda)$ obtained using the numerical solution for $\theta(\lambda)$ is plotted in Fig. 6 for two different initial values of $r_{0}$. The approximate expression in (68) is in good agreement with the numerical solution for both the initial conditions. Furthermore, from this solution for $r(\lambda)$, we once again see that one can reach any arbitrary value of $r$, for a finite value of the parameter $\lambda$, following this geodesic. Therefore, any point arbitrarily far away from the integrable phase can be reached with a finite distance, i.e., a finite value of the ’complexity’.
@@
Before concluding this section, we note that here we have performed an analysis of different geometric quantities for the generalised version of the conventional RP model with $N=2$. While it would be very much desirable to perform a similar analysis for a generic $N$, analytic expressions for the QMT components for an arbitrary value of $N$ are not known. In [44], these components were obtained numerically, and it was shown that the metric components follow different scaling properties in different phases of this model. E.g., when $r\ll r^{*}=1/\sqrt{N}$ (corresponding to the localised phase of the model), the metric components follow a $1/r$ scaling, while in the ergodic phase, having $r\gg 1$, the metric components approach those of the GUE random matrix expressions in (27). Finally, in the non-ergodic extended phase ($r^{*}\ll r\ll 1$), the components $\tilde{g}_{rr}$ and $\tilde{g}_{\phi\phi}/r^{2}$ goes to constants scaling as $N$.
@@ VI Geometry of quantum states for Gaussian $\beta$-ensembles
In this section and the next, we consider a generalisation of the integrability-breaking Hamiltonian considered in the previous section, such that the Dyson index ($\beta$) can take a range of real positive values, not just those of classical Gaussian ensembles. Specifically, in this section, we consider the tridiagonal matrix representation of the so-called Gaussian $\beta$-ensembles, found in [42]. A Hamiltonian belonging to this representation of the ensemble takes a real symmetric tridiagonal $N\times N$ matrix form whose diagonal elements (denoted as $a_{n}$, $n=0,\cdots N-1$) are are independent Gaussian random variables with mean 0, and variance as that of the ensemble, while the sub-diagonal elements (which we call $b_{n}$, with $n=1,\cdots N-1$) are independently distributed random variables with the following distribution
@@
A natural property of the ensemble of Hamiltonians with general $\beta$ is that, by controlling the value of the Dyson index, one can control the strength of the repulsion between the energy levels. In the limit $\beta\rightarrow 0$, the energy levels can be arbitrarily close to each other without any level repulsion, and therefore the level spacing statistics is Poissonian, associated with integrable systems. Increasing the value of $\beta$ from this value, one gradually gets ensembles with different degrees of level repulsion, including those of the classical Gaussian ensembles (with $\beta=1,2,4$), as well as ensembles having intermediate values of $\beta$. Therefore, the Gaussian $\beta$-ensembles can be thought to be showing a chaos-to-integrability transition as $\beta$ is decreased below the GUE value ($\beta=1$) towards zero. Similar to the RP model, apart from the chaos to integrability transition, the $\beta$-ensembles also show Anderson localisation transition for a specific value of the Dyson index. Rewriting $\beta=N^{\gamma}$ for some real parameter $\gamma$, it has been shown recently that $\beta$-ensembles have three different phases, respectively, an ergodic phase ($\gamma\leq 0$), a nonergodic extended phase (for $0<\gamma<1$) and a localised phase when $\gamma\geq 1$ [79].
@@
For these ensemble with generic Dyson index, one can follow two different procedures to move to the integrable phase: 1. first, as was the case in section III, we can add a term proportional to random matrix drawn from an ensemble with a fixed general value of the Dyson index $\beta>1$ to a diagonal matrix having independent random variables as the entry. As one takes the proportionality constant (denoted as $r$) to zero, the total Hamiltonian makes a transition from chaotic to the integrable phase as $r\rightarrow 0$. 2. Secondly, we can consider a Hamiltonian which is a sum of two terms, each of which is taken from these generalised $\beta$-ensembles and has, in general, different Dyson indices ($\beta_{1}$ and $\beta_{2}$). Then, as one takes either or both of these $\beta_{i}$ close to zero and takes $r\rightarrow 0$, the total Hamiltonian again makes a transition from chaotic to an integrable phase.
In this section, we follow the first option, i.e., we take a rank-2 matrix sampled from the tridiagonal Gaussian $\beta$-ensembles, add it to a diagonal integrable Hamiltonian, and derive the corresponding ensemble-averaged FS (which is the $rr$ component of the line element in (58)). [^16] One reason for following this procedure is that, as we shall see, the integral for the FS does not converge for the tridiagonal Hamiltonian when $\beta<1$. We first compute this for the case with $\beta=2$, and show that it matches with the expression in (58), and subsequently generalise the results to obtain the ensemble-averaged FS for a generic value of $\beta>1$. [^17] In the next section VII, we have provided the computation of the FS of a class of random matrices with general Dyson index, which, in contrast to the tridiagonal form used in this section, is rotationally invariant. Also in Appendix E we present a computation of the FS for a random matrix ensemble with a generic Dyson index $\beta$.